your faimly likes to go to baseball games. at one game your faimly bought 5 soft drinks and 5 hot dogs for $22.25. at the next game your faimly attended the bought 4 soft drinks and 3 hot dogs for $14.50. what is the cost of one soft drink and one hot dog?
let s = the cost of one soft drink
let h = the cost of one hot dog
a. write a system of equations modeling the situation described above.
b. solve the system for the cost of one soft drnik + one hot dog.

Answer by ankor@dixie-net.com(22740)

Your family bought 5 soft drinks and 5 hot dogs for $22.25. At the next game your family attended, they bought 4 soft drinks, and 3 hot dogs for $14.50. What is the cost of one soft drink and one hot dog?
:
let s = the cost of one soft drink
let h = the cost of one hot dog
:
a. write a system of equations modeling the situation described above.
5s + 5h = 22.25
4s + 3h = 14.50
:
b. solve the system for the cost of one soft drink + one hot dog.
Multiply the 1st equation by 3 and the 2nd equation by 5, resulting in:
:
15s + 15h = 66.75
20s + 15h = 72.50
--------------------subtracting eliminates h:
-5s + 0h = -5.75
-5s = -5.75
s = -5.75/5
s = $1.15 ea for soft drinks
:
Find h using the 1st equation, substitute 1.15 for s
5s + 5h = 22.25
5(1.15) + 5h = 22.25
5.75 + 5h = 22.25
5h = 22.25 - 5.75
5h = 16.5
h = 16.5/5
h = $3.30 ea for hot-dogs
:
:
Check solution in 2nd equation:
4(1.15) + 3(3.30) =
4.60 + 9.90 = 14.50
:
They want the cost of 1 hot-dog and 1 drink, I'll let you figure that out
Does this make sense to you? Any questions?